80cm3 mixture of propane and hydrogen was mixed with excess oxygen and exploded. The product after cooling to the original room temperature were treated with excess potassium hydro oxide solution when the volume decrease by 60cm³. Calculate the composition of the original mixture


1
Expert's answer
2021-03-15T09:10:35-0400

To get the volume of O2 :


=(80-60)=20cm3


The composition of the original mixture =(80+20)cm3


=100cm3